What is the pH of a 40.0 mL solution that is 0.17 M in CN and 0.25 M in HCN? The K, for HCN is 4.9 x 10-9 pH=
Answer #1

in accordance with Henderson Hasselbalch equation

pH = pKa + log ([salt] /[acid])

pH = pKa +log ([CN-] /[HCN]) ...... (1)

given Ka =4.9*10^(-9)

[CN-] =0.17M

[HCN] =0.25M

pKa = - log Ka = - log(4.9*10^(-9)) =8.31

[CN-]/[HCN] = 0.17/0.25 =0.68

put values in (1)

pH =8.31 + log 0.68 = 8.14

thus we calculated pH using Henderson Hasselbalch equation.